OpenStudy (anonymous):

Given that line segment QT is a median of triangle QRS, find RS. A.2 B.4 C.9 D.18 I got 2 for this answer..just wanna check to make sure.

OpenStudy (anonymous):



OpenStudy (anonymous):

i tried doing this one 2 different ways. both were wrong answers :/

hero (hero):

T is the midpoint of RS, so RT = TS Therefore: 3x + 3 = 7x - 5 Solving for x, you get: 3 + 5 = (7 - 3)x 8 = 4x 8/4 = x 2 = x But we still have to find RS

OpenStudy (anonymous):

i got the x=2 part but i didnt kno where to go from there.

hero (hero):

RT = 3x + 3 = 3(2) + 3 = 6 + 3 = 9 RT = TS so TS = 9 as well 9 + 9 = 18 (again)

hero (hero):

So RS = 18

OpenStudy (anonymous):

ohhh, yeah thats where i was stuck. forgot to replace x with 2 .

OpenStudy (anonymous):

weird that i got the same answer for a different problem.. :/ but thanks for helping again! (: